Which of the following matrix (or given expression) results in a non-singular matrix


1) 2 0 0

0 SINà ½SINà

0 -1 -COSà


2) A v B where A = 1 2 -1

3 1 0

-2 -4 2


and B = 1 0 -1

3 1 0

-2 -4 2


3) 2 0 0

3 -4 0

0 -1 0


4) 2 0 0

3 -4 4

0 1 -1


1
Expert's answer
2021-04-29T16:16:59-0400

.A square matrix is non singular iff its determinant is non zero.

Consider matrices A and B in 2

det A="\begin{vmatrix}
 1 & 2 & -1 \\
 3 & 1 & 0\\
 -2 & -4 & 2
\end{vmatrix}" = 1·1·2 + 2·0·(-2) + (-1)·3·(-4) - (-1)·1·(-2) - 1·0·(-4) - 2·3·2 = 2 + 0 + 12 - 2 - 0 - 12 = 0

det В = "\begin{vmatrix}
 1 & 0 & -1 \\
 3 & 1 & 0\\
 -2 & -4 & 2
\end{vmatrix}" = ·(-1)·2 + 0·0·(-2) + (-2)·3·(-4) - (-2)·(-1)·(-2) - 1·0·(-4) - 0·3·2 = -2 + 0 + 24 + 4 - 0 - 0 = 26

Since det A=0 and det B=26 then det(A"\lor" B)=26

so matrix A"\lor" B is non-singular